Consider the following.
g(y) = 7 − 2y
Evaluate the function at each specified value of the independent variable and simplify. (If the function is undefined at the specified value, enter UNDEFINED.)
(a) g(0)

(b) g(7/2)

(c) g(s + 2)

Answer by RAY100(1637)

g(y) = 7-2(y)
.
g(0) = 7 - 2(0)=7,,,,,subst 0 for y
.
g(7/2) = 7-2(7/2) = 7-7=0,,,,subst (7/2) for y
.
g(s+2) =7-2(s+2) =7-2s-2(2) = 7-2s -4 = 3-2s ,,,,,subst (s+2) for y